OpenStudy (anonymous):

The perimeter of a rectangle is 54cm. One side is x cm long, and the other is 6cm longer. a) form an equation involving x b) Solve the equation, and write down the length of each of the sides.

OpenStudy (mayankdevnani):

\[\huge perimeter~of~rectangle=2(l+b)\]

OpenStudy (mayankdevnani):

if length = 6 cm and breadth = x cm

OpenStudy (mayankdevnani):

so, it becomes: \[\huge perimeter~of~rectangle=2(6+x)\] \[\huge perimeter=12+2x\]

OpenStudy (jack1):

"The perimeter of a rectangle is 54cm. One side is x cm long, and the other is 6cm LONGER...." @mayankdevnani i think the equations should actually be : B = x L = x+6 P = 54 so if Perimeter = 2(L + W) = 54 \[P = 2 ((x+6) + x)\] so \[P = 4x+12\]

OpenStudy (mayankdevnani):

@Jack1 but in question it says that it is 6 cm longer not 6 cm longer than breadth

OpenStudy (mayankdevnani):

so , \[\huge 54=2(6+x)\] \[\huge 54=12+2x\] \[\huge 2x=42\] \[\huge x=21\] so breadth =21 cm

OpenStudy (jack1):

sorry dude, i think the grammar of the statement implies that it is longer than breadth, even if it doesn't state it outright

OpenStudy (mayankdevnani):

yeah!!! no problem its human misunderstanding

OpenStudy (mayankdevnani):

but i will reward you because you try it

OpenStudy (jack1):

human... you... i'm pretty sure you're a robot, you're always on OS ;P

OpenStudy (mayankdevnani):

whether it is wrong or right

OpenStudy (jack1):

lol, ur such a sore loser sometimes jk ;D

OpenStudy (mayankdevnani):

always???? no way...

OpenStudy (jack1):

;D

OpenStudy (mayankdevnani):

@Jack1 i came on OS regularly because my summer vacations is going on and my summer vacations end up in 2 days and i can't open OS regularly..i think i open it once in a week

OpenStudy (mayankdevnani):

okk.. @Jack1

OpenStudy (jack1):

ah ha ha ha ... dude i was just messin around, you're taking it way too seriously .... ...(unless you really are a maths solving robot...) ;D We love u here dude, u brighten up our chat hey!

OpenStudy (mayankdevnani):

lol

OpenStudy (mayankdevnani):

i am not math solving robot...i am poor in studies

OpenStudy (jack1):

but rich in friends

OpenStudy (jack1):

k we've taken over this guys post for long enough, back to the chat?

OpenStudy (mayankdevnani):

okk...

OpenStudy (anonymous):

p = 2x + 12 you solve 2x + 12 = 54 2x = 54 - 12 2x = 42 x = 21 p = 54